OpenStudy (anonymous):

if 48 square inches are substacted for the area of a square table the resut would be equal to twice the peremiter. what is the legth of the side of the table

OpenStudy (anonymous):

Let the side of the square table = x inch then its perimeter = 4*side = 4*x = 4x inch area = side*side = x*x = x^2 sq.inch ATQ, x^2 - 48 = 2(4x) x^2 - 48 = 8x x^2 - 8x - 48 = 0 x^2 - 12x + 4x - 48 = 0 x(x - 12) + 4(x - 12) = 0 (x + 4)(x - 12) = 0 so if x + 4 = 0 then x = -4 if x - 12 = 0 then x = 12 since length can never be negative, hence x = 12 is the acceptable value Hence the length of the side of the square table is 12 inch